OpenStudy (anonymous):

Use the formulas for lowering powers to rewrite the expression in terms of the first power of cosine. -cos(x)sin^2(x) -cos^7(x)

OpenStudy (anonymous):

Use the double angle formula \[ \cos^2(x) =\frac {1+\cos(2x)}2\\ \sin^2(x) =\frac {1-\cos(2x)}2\\ \]

OpenStudy (anonymous):

yes I have used the formulas. I used cos x sin^2x and have gotten it wrong?? please help :( \[(\cos(x)-\cos^2(2x))/2 \]

OpenStudy (anonymous):

\[ -\cos(x)\sin^2(x)=\frac {-\cos(x)+\cos(x)\cos(2x)}2\\ \]

OpenStudy (anonymous):

isn't cos(x)cos(2x) the same thing??

OpenStudy (anonymous):

--- as cos^2x?